# Stable infinite-temperature eigenstates in SU(2)-symmetric nonintegrable models

### v1

@@ Abstract
Introduction.— The concept of nonintegrability is closely associated with the thermalization of all eigenstates of a many-body quantum system, alongside their complex chaotic behavior [1, 2, 3, 4]. At long times, the average properties of such systems can be described using the microcanonical ensemble of statistical mechanics, effectively erasing all information about initial conditions to local observables [1, 5, 6]. Despite this statistical description, nonintegrable models remain analytically intractable [7, 8, 9]: unlike integrable systems, they lack an extensive set of local integrals of motion [10], allowing primarily for numerical treatments.
However, certain nonintegrable models have recently been found to defy such statistical description. Examples include many-body quantum scars [11, 12, 13, 14, 15, 16, 17, 18, 19], which are atypical states within the spectrum which violate the eigenstate thermalization hypothesis and sometimes can be constructed exactly. Often these are characterised by low entanglement, but sometimes occur with volume-law entanglement [20, 21, 22, 23]. These out-of-equilibrium states are gaining interest due to their potential applications for quantum computing architectures [24].
@@
This raises questions about whether zero-energy subspaces in other models might demonstrate analogous fascinating characteristics. Of particular interest would be an exploration of models with complex internal symmetries that have not been studied before. In this Letter, we investigate a class of nonintegrable bond-staggered (alternating) spin chains that possesses a non-Abelian internal symmetry. These models have attracted both theoretical [39, 40, 41] and experimental [42, 43] interest due to their connection to Haldane spin chains and resemblance to the Su-Schrieffer-Heeger model. Here, we employ character theory to obtain nullspace degeneracies within each symmetry sector, which we find to be exponentially diverging with the system size. Through symmetry considerations, we are able to construct a non-trivial exact zero-energy basis, serving as a few-body scarred subspace. This fixed-separation basis possesses low entanglement and is remarkably stable to experimental perturbations. These results push our current understanding of the dynamical stability of special states in non-Abelian chaotic models.
@@
where $\mathbf{S}_{i}=(S^{x}_{i},S^{y}_{i},S^{z}_{i})$ is the vector of spin-half operators, and $V$ is any perturbation that preserves the symmetries of the model, as defined below. The presence of the perturbation term makes our results generic while allowing for investigations of stability against external disruptions. The second equality allows us to interpret the Hamiltonian as a collection of SWAP gates, which will prove useful in subsequent analysis. The model is endowed with a non-Abelian SU(2) total spin symmetry and the total spin conservation in any direction [U(1) symmetry]; without a loss of generality, we consider conservation of $S^{z}_{\text{tot}}=\sum_{i}S^{z}_{i}$. This forms the internal symmetry group $\mathcal{G}$. The Hamiltonian (1a) anticommutes with the translation operator $\tau$ and commutes with translation by two sites, $\tau^{2}$, and with bond-reflection symmetry, $\sigma$. Together, these form the spatial symmetry group of the model, the dihedral group $D_{2L}$, illustrated in Fig. 1(a). We term the group generated by $\tau$ and $\sigma$ the anti-symmetry group $\mathcal{A}$.
Previous research [44, 45] has implied that staggering the interaction term is responsible for breaking the integrability of the Heisenberg chain. We show this explicitly through level statistics for the periodic model with $V=0$ by resolving all its symmetries. We write the Hamiltonian in fixed momentum basis [46], which resolves $S^{z}_{\text{tot}}$ and $\tau^{2}$. $\mathbf{S}^{2}$ is resolved by comparing the spectra of two consecutive values of $S^{z}_{\text{tot}}$, while $\sigma$ is diagonalized directly. We then plot the distribution of the ratio of consecutive level spacings $r$ [47] in Fig. 1(b), which we find to match the prediction for the Gaussian orthogonal ensemble (GOE) of random matrix theory [48]. This implies the model is chaotic, strongly suggesting its nonintegrability. Moreover, the distribution exhibits level repulsion at $r=0$, while if the model had some remaining symmetries or was integrable, repulsion would be absent and the histogram would resemble the Poisson distribution.
@@
Zero-energy states enforced by the symmetry.— Through symmetry considerations, we calculate the lower bound on the nullspace dimension of the Hamiltonian and compare the result with exact diagonalization. The intuition for why the symmetry can inform about ZED is as follows. Consider any Hamiltonian $H$ that anticommutes with some unitary operator $A$, thus implying a symmetric spectrum. Suppose that $A$ has only real eigenvalues, ±$\pm 1$, with multiplicities $m_{+}$ and $m_{-}$. In the eigenbasis of $A$, the Hamiltonian takes a block off-diagonal form, consisting of blocks of sizes $(m_{+}\times m_{-})$. Consequently, $H^{2}=H^{\dagger}H$ becomes block-diagonal, separating into two sectors of $A$, with dimensions $(m_{+}\times m_{+})$ and $(m_{-}\times m_{-})$, both necessarily of the same rank. If there is an imbalance between $m_{+}$ and $m_{-}$, then the nullspace of $H^{2}$ (and consequently the nullspace of $H$) is at least $|m_{+}-m_{-}|$-dimensional. Note that ZED may be larger if both sectors contain zero-energy states. However, if the model is chaotic, the lower bound is expected to be saturated.
@@
where $m_{\kappa,s}$ are the multiplicities of the antisymmetry representation labeled by $(\kappa,s)$, with the $s$ label suppressed in the second sum as it is forced by $\kappa$. This is enhanced by internal symmetries which cause the same computation to be repeated in each $\mathcal{G}$ sector separately.
We can calculate these symmetry-resolved multiplicities using character theory [49]. Specifically, taking the character function $\chi_{\mathcal{H}}$ for the representation on the full Hilbert space $\mathcal{H}$, we can use the Schur orthogonality relations for the natural inner product to extract the multiplicity of each irreducible representation within the full representation. Due to the product structure of the group we are interested in, we can write the full character as a function of $g\in\mathcal{A}$ and $\theta\in\mathcal{G}$. This allows the inner product to be taken in two steps, first over the spatial part to produce a kind of generalized Schur function, $F_{\kappa,s}(\theta)=\langle\chi_{\kappa,s}(g),\chi_{\mathcal{H}}(g,\theta)\rangle_{\mathcal{A}}$. Note that $\chi_{\kappa,s}(\sigma^{\nu}\tau^{\mu})=d_{\kappa}s^{\nu}\cos(2\pi\kappa\mu/L)$, where $d_{\kappa}\in\{1,2\}$ is the dimensionality of irrep $\kappa$  [^1] See Appendix A in the Supplemental Material for more details about character theory, tables of zero-energy degeneracies, and discussion of extra zero-energy states.. Second, over the internal symmetry group to extract the multiplicity of an individual representation,
@@
where $S$ labels the total spin sector.
Now, using the cycle decomposition of $g$, viewed as a permutation of sites,
@@
where $c_{l}(g)$ is the $l$-cycle index of $g$ which evaluates to the number of cycles of length $l$ in $g$. The final trace appearing in Eq. (4) is over a system of size $l$ (Hilbert space $\mathcal{H}_{l}$) and we call this the $l$-cycle character function. Analogously to its use for the single-site translation, we use $\tau$ here to refer to the generator of the cyclic group within each cycle.
@@
At this point, we specialize to the main internal symmetry of interest, SU(2). The character function for the irreducible representation of spin-($l/2$) is $\chi_{l/2}^{\text{SU(2)}}(\theta)=\sin[(l+1)\theta]/\sin\theta$ [49], and this forms a basis for the space of character functions. Pointwise operations on these produce a ring, known variously as the character ring or representation ring, and are equivalent to the fusion rules of SU(2) [51]. This equivalence takes the tensor product of representations to pointwise-multiplication of characters and direct sum to addition, producing equations such as $\chi_{1/2}\chi_{1/2}=\chi_{0}+\chi_{1}$ from the fusion rules of SU(2). The character function for the $l$-cycle evaluates to
$\mathrm{tr}_{\mathcal{H}_{l}}\!\left(\tau\theta\right)=2\cos l\theta=\chi_{l/2}(\theta)-\chi_{l/2-1}(\theta).$ (5)
Finally, we perform polynomial-complexity arithmetic in this ring, and combine Eqs. (25) to compute the zero-energy degeneracies in any sector.
We compare the number of zero-energy eigenstates predicted by the character theory with the exact diagonalization results for $V=0$ in Table 1 and find an almost exact match, except for a few extra zero-energy states for small system sizes [50]. This additional degeneracy is most likely accidental, occurring due to small symmetry sector size, rather than from an unresolved symmetry, for which we see no evidence. The degeneracy can be lifted with the addition of a perturbation, such as $V=\lambda\sum_{i}(-1)^{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+3}$ for a generic perturbation strength $\lambda$, while respecting the symmetries. Another interesting finding is a correspondence between multiplicities in sectors $\{S=L/2-2n,k,s\}$ and $\{S-1,\pi-k,-s\}$, which we observe in character theory results across all available system sizes (up to $L=42$). We note however that this correspondence is a mathematical property coming from the arithmetics of the cycles in Eq. (5), and does not imply the existence of a simple operator that maps states between the two sectors.
@@
Each orbit contains a so-called 00 representation denoted $\chi_{0}$ which sets $\chi(\tau)=+1$ and gives an orbit-dependent sign for $\chi(\sigma)$. Every orbit, except for the orbit with separation $L/2$, has a so-called $\pi$ representation denoted $\chi_{\pi}$ which sets $\chi(\tau)=-1$ and again gives an orbit dependent sign for $\chi(\sigma)$. The reflection-parity signs will be irrelevant for everything that follows, justifying suppressing them from the notation.
@@
Entanglement and stability of the fixed-separation basis.— Given their analytical tractability, we inquire whether these zero-energy states are endowed with interesting physical properties, such as being few-body quantum scars. We discover that the states with fixed magnon separation exhibit an area-law von Neumann entanglement entropy scaling [see Fig. 3(a)]  [^3] See Appendix C in the Supplemental Material for detailed calculations of the von Neumann energy of fixed-separation states., which tends to $\ln 2$ in the infinite volume limit. Conversely, if the separation scales linearly with the system size, the entropy behaves as expected for a generic state in a two-magnon sector, i.e., $\sim\ln L$.
@@
where $V$ is the perturbation in question and $\lambda$ is its strength. We use representative examples of even/odd separation states to be those of $x=2$ and $x=3$, respectively. As shown in Fig. 3(b), we find that the basis is remarkably stable: states of even separations are only unstable to the staggered field and $\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+2}$, while odd-separation states are unstable to Heisenberg-type interaction terms. This implies that one can form a stable fixed-separation basis of zero-energy states, which can be used to encode some initial quantum information. This information is then preserved for a long period of time if the strength of the destabilizing perturbations is small. Interestingly, the Loschmidt echo for the staggered field perturbation shows an unusually strong revival at $t\approx 1,\lambda=1.0$, implying that one can recover the initial state even if the additional term generally destroys the coherence of the state.
@@ Acknowledgements.
Acknowledgments.— A. P., B. M., and M. S. were funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 853368). C. J. T. is supported by an EPSRC fellowship (Grant Ref. EP/W005743/1). The authors acknowledge the use of the UCL High Performance Computing Facilities (Myriad and Kathleen), and associated support services, in the completion of this work. R. M. and H. J. C. acknowledge funding from NSF DMR 2046570 and the National High Magnetic Field Laboratory. The National High Magnetic Field Laboratory is supported by the National Science Foundation through NSF/DMR-1644779 and DMR-2128556 and the state of Florida.
@@ I.1 Character theory details
Let us start by explicitly writing the representations for the dihedral group, which is the anti-symmetry group $\mathcal{A}$ of our system (see Fig. 1 in the main text). It contains one-dimensional irreps labeled by $\kappa=0,L/2$ and $s=\pm 1$ with representation
@@
The characters are defined as traces of representation matrices, and hence it is easy to see that for 1D irreps $\chi_{\kappa,s}(\sigma^{\nu}\tau^{\mu})=s^{\nu}\omega_{\kappa}^{\mu}$, while for 2D irreps $\chi_{\kappa}(\sigma^{\nu}\tau^{\mu})=2\cos(2\pi\kappa\mu/L)$. We can write this succinctly as $\chi_{\kappa,s}(\sigma^{\nu}\tau^{\mu})=d_{\kappa}s^{\nu}\cos(2\pi\kappa\mu/L)$, where $d_{\kappa}\in\{1,2\}$ is the dimensionality of irrep $\kappa$ and we assume that for 2D irreps we take $s=+1$.
Now, we turn to the discussion of characters for the internal symmetry of the model, SU(2). Consider the character function for the system of size $l$ (Hilbert space $\mathcal{H}_{l}$) from Eq. (4) in the main text. The trace can be expressed in some product orthonormal basis
@@
$=\sum_{\alpha_{1},\dots,\alpha_{l}}\bra{\alpha_{1}}\rho(\theta)\ket{\alpha_{2}}\cdots\bra{\alpha_{l}}\rho(\theta)\ket{\alpha_{1}}$ (S.3b)
$=\sum_{\alpha_{1}}\bra{\alpha_{1}}\rho(\theta)^{l}\ket{\alpha_{1}}=\mathrm{tr}[\rho(\theta)^{l}],$ (S.3c)
where $\rho(\theta)$ is the matrix representation of the SU(2) group, and |$\ket{\alpha_{i}}$ are the basis elements. Now we can note that any spin-1/2 SU(2) matrix can be transformed into a diagonal form $\rho(\theta)=\operatorname{diag}(e^{i\theta},e^{-i\theta})$, giving us an easy way to compute the trace and characters,
@@
×$\quad\times\sum_{\mu=0}^{l-1}\sum_{\nu=0}^{1}d_{k}s^{\nu}\cos\frac{2\pi\kappa\mu}{L}\prod_{l}(2\cos\theta l)^{c_{l}(g)}.$
This can be evaluated explicitly after first performing polynomial-complexity calculations to obtain cycles and their lengths $c_{l}(g)$.

### v2

@@ Abstract
Introduction.— The concept of nonintegrability is closely associated with the thermalization of all eigenstates of a many-body quantum system, alongside their complex chaotic behavior [1, 2, 3, 4]. At long times, the average properties of such systems can be described using the microcanonical ensemble of statistical mechanics, effectively erasing all information about initial conditions to local observables [1, 5, 6]. Despite this statistical description, nonintegrable models remain analytically intractable [7, 8, 9]: unlike integrable systems, they lack an extensive set of local integrals of motion [10], leaving numerical treatments our primary source of understanding.
However, certain nonintegrable models have recently been found to defy such a statistical description. Examples include systems with many-body quantum scars [11, 12, 13, 14, 15, 16, 17, 18, 19], which are atypical states within the spectrum which violate the eigenstate thermalization hypothesis and sometimes can be constructed exactly. Often these are characterized by low entanglement, but sometimes occur with volume-law entanglement [20, 21, 22, 23]. These out-of-equilibrium states are gaining interest due to their potential applications for quantum computing architectures [24].
@@
This raises questions about whether zero-energy subspaces in other models [39] might demonstrate analogous fascinating characteristics. Of particular interest would be an exploration of models with complex internal symmetries that have not been studied before. In this Letter, we investigate a class of nonintegrable bond-staggered (alternating) spin chains that possesses a non-Abelian internal symmetry. These models have attracted both theoretical [40, 41, 42] and experimental [43, 44] interest due to their connection to Haldane spin chains and resemblance to the Su-Schrieffer-Heeger model. Here, we employ character theory to obtain nullspace degeneracies within each symmetry sector, which we find to be scaling exponentially with the system size. Through symmetry considerations, we are able to construct a non-trivial exact zero-energy basis, serving as a few-body scarred subspace. This fixed-separation basis possesses low entanglement and is remarkably stable to experimental perturbations. These results push our current understanding of the dynamical stability of special states in non-Abelian chaotic models.
@@
where $\mathbf{S}_{i}=(S^{x}_{i},S^{y}_{i},S^{z}_{i})$ is the vector of spin-half operators, and $V$ is any perturbation that preserves the symmetries of the model, as defined below. The presence of the perturbation term makes our results generic while allowing for investigations of stability against external disruptions. The second equality allows us to interpret the Hamiltonian as a collection of SWAP gates, which will prove useful in subsequent analysis. The model is endowed with a non-Abelian SU(2) total spin symmetry and the total spin conservation in any direction [U(1) symmetry]; without a loss of generality, we consider conservation of $S^{z}_{\text{tot}}=\sum_{i}S^{z}_{i}$. This forms the internal symmetry group $\mathcal{G}$. The Hamiltonian (1a) anticommutes with the translation operator $\tau$ and commutes with translation by two sites, $\tau^{2}$, and with bond-reflection symmetry, $\sigma$. Together, these form the spatial symmetry group of the model, the dihedral group $D_{L}$, illustrated in Fig. 1(a). We term the $D_{2L}$ group generated by $\tau$ and $\sigma$ the anti-symmetry group $\mathcal{A}$.
Previous research [45, 46] has implied that staggering the interaction term is responsible for breaking the integrability of the Heisenberg chain. We show this explicitly through level statistics for the periodic model with $V=0$ by resolving all its symmetries. We write the Hamiltonian in the fixed momentum magnon basis [47], which resolves $S^{z}_{\text{tot}}$ and $\tau^{2}$. $\mathbf{S}^{2}$ is resolved by comparing the spectra of two consecutive values of $S^{z}_{\text{tot}}$, while $\sigma$ is diagonalized directly. We then plot the distribution of the ratio of consecutive level spacings $r$ [48] in Fig. 1(b), which we find to match the prediction for the Gaussian orthogonal ensemble (GOE) of random matrix theory [49]. This implies the model is chaotic, strongly suggesting its nonintegrability. Moreover, the distribution exhibits level repulsion at $r=0$, while if the model had some remaining symmetries or was integrable, repulsion would be absent and the histogram would resemble the Poisson distribution.
@@
Zero-energy states enforced by the symmetry.— Through symmetry considerations, we calculate the lower bound on the nullspace dimension of the Hamiltonian and compare the result with exact diagonalization. The intuition for why the symmetry can inform about ZED is as follows. Consider any Hamiltonian $H$ that anticommutes with some unitary operator $A$, thus implying a symmetric spectrum ($E\mapsto-E$). Suppose that $A$ has only real eigenvalues, ±$\pm 1$, with multiplicities $m_{+}$ and $m_{-}$. In the eigenbasis of $A$, the Hamiltonian takes a block off-diagonal form, consisting of blocks of sizes $(m_{+}\times m_{-})$. Consequently, $H^{2}=H^{\dagger}H$ becomes block-diagonal, separating into two sectors of $A$, with dimensions $(m_{+}\times m_{+})$ and $(m_{-}\times m_{-})$, both necessarily of the same rank. If there is an imbalance between $m_{+}$ and $m_{-}$, then the nullspace of $H^{2}$ (and consequently the nullspace of $H$) is at least $|m_{+}-m_{-}|$-dimensional [50]. Note that ZED may be larger if both sectors contain zero-energy states. However, if the model is chaotic, the lower bound is expected to be saturated.
@@
where $m_{\kappa,s}$ are the multiplicities of the antisymmetry representation labeled by $(\kappa,s)$, with the $s$ label suppressed in the second sum as it is forced by $\kappa$. This is enhanced by internal symmetries, which cause the computation to be performed in each $\mathcal{G}$ sector separately.
We can calculate these symmetry-resolved multiplicities using character theory [51, 52]. The character function $\chi_{\rho}(g)=\operatorname{tr}(\rho(g))$ is a function that identifies a representation $\rho$. Specifically, we take the character function for the representation on the full Hilbert space $\mathcal{H}$ and we use the Schur orthogonality relations for the natural inner product: the multiplicity of each irreducible representation within that full representation is the inner product of the irrep character and the character of the full representation. Due to the product structure of the group we are interested in, we can write this full character as a function of $g\in\mathcal{A}$ and $\theta\in\mathcal{G}$. This allows the inner product to be taken in two steps, first over the spatial part to produce $F_{\kappa,s}(\theta)=\langle\chi^{\mathcal{A}}_{\kappa,s}(g),\chi_{\mathcal{H}}(g,\theta)\rangle_{\mathcal{A}}$. Note that $\chi^{\mathcal{A}}_{\kappa,s}(\sigma^{\nu}\tau^{\mu})=d_{\kappa}s^{\nu}\cos(2\pi\kappa\mu/L)$, where $d_{\kappa}\in\{1,2\}$ is the dimensionality of irrep $\kappa$  [^1] See Appendix A in the Supplemental Material for more details about character theory, tables of zero-energy degeneracies, and discussion of extra zero-energy states.. Second, over the internal symmetry group to extract the multiplicity of an individual representation,
@@
where $S$ labels the total spin sector. Note that one does not need to resolve $S^{z}_{\text{tot}}$, as it only acts inside each SU(2) representation.
Now, viewing $g$ as a permutation of sites, we can use its decomposition into cycles of sites,
@@
where $c_{l}(g)$ is the $l$-cycle index of $g$ which counts the number of cycles of length $l$ in $g$ [53]. The final trace appearing in Eq. (4) is over a system of size $l$ (Hilbert space $\mathcal{H}_{l}$) and we call this the $l$-cycle character function. Analogously to its use for the single-site translation, we use $\tau$ here to refer to the generator of the cyclic group within each cycle.
@@
At this point, we specialize to the main internal symmetry of interest, SU(2). The character function for the irreducible representation of spin-$S$ is $\chi_{S}^{\text{SU(2)}}(\theta)=\sin[(2S+1)\theta]/\sin\theta$ [51], and this forms a basis for the space of character functions. Pointwise operations on these produce a ring, known variously as the character ring or representation ring, and are equivalent to the fusion rules of SU(2) [54]. This equivalence takes the tensor product of representations to pointwise-multiplication of characters and direct sum to addition, producing equations such as $\chi^{\text{SU(2)}}_{1/2}\chi^{\text{SU(2)}}_{1/2}=\chi^{\text{SU(2)}}_{0}+\chi^{\text{SU(2)}}_{1}$ from the fusion rules of SU(2). The character function for the $l$-cycle evaluates to
$\mathrm{tr}_{\mathcal{H}_{l}}\!\left(\tau\theta\right)=2\cos l\theta=\chi^{\text{SU(2)}}_{l/2}(\theta)-\chi^{\text{SU(2)}}_{l/2-1}(\theta).$ (5)
Finally, we perform polynomial-complexity arithmetic in this ring [53], and combine Eqs. (25) to compute the zero-energy degeneracies in any sector.
We compare the number of zero-energy eigenstates predicted by the character theory with the exact diagonalization results for $V=0$ in Table 1 and find an almost exact match, except for a few extra zero-energy states for small system sizes [53]. This additional degeneracy is most likely accidental, occurring due to small symmetry sector size, rather than from an unresolved symmetry, for which we see no evidence. The degeneracy can be lifted with the addition of a perturbation, such as $V=\lambda\sum_{i}(-1)^{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+3}$ for a generic perturbation strength $\lambda$, while respecting the symmetries. Another interesting finding is a correspondence between multiplicities in sectors $\{S=L/2-2n,k,s\}$ and $\{S-1,\pi-k,-s\}$, which we observe in character theory results across all available system sizes (up to $L=42$). We note however that this correspondence is a mathematical property coming from the arithmetic of the cycles in Eq. (5), and does not imply the existence of a simple operator that maps states between the two sectors.
@@
Each orbit contains a so-called 00 representation denoted $\chi_{0}$ which sets $\chi(\tau)=+1$ and gives an orbit-dependent sign for $\chi(\sigma)$. Every orbit, except for the orbit with separation $L/2$, has a so-called $\pi$ representation denoted $\chi_{\pi}$ which sets $\chi(\tau)=-1$ and again gives an orbit dependent sign for $\chi(\sigma)$. The reflection-parity signs are irrelevant for the fixed-separation states, justifying suppressing them from the notation.
@@
Entanglement and stability of the fixed-separation basis.— Given their analytical tractability, we inquire whether these zero-energy states are endowed with interesting physical properties, such as being interacting quantum scars. We discover that the states with fixed magnon separation exhibit an area-law von Neumann entanglement entropy scaling [see Fig. 3(a)]  [^3] See Appendix C in the Supplemental Material for detailed calculations of the von Neumann energy of fixed-separation states., whose value approaches $\ln 2$ in the infinite volume limit. Conversely, if the separation scales linearly with the system size, the entropy behaves as expected for a generic state in a two-magnon sector, i.e., $\sim\ln L$.
@@
where $V$ is the perturbation in question and $\lambda$ is its strength. We use representative examples of even/odd separation states to be those of $x=2$ and $x=3$, respectively. In each class the qualitative behavior is the same. As shown in Fig. 3(b), we find that the basis is remarkably stable: states of even separations are only unstable to the staggered field and $\sum_{i}\mathbf{S}_{i}\cdot\mathbf{S}_{i+2}$, while odd-separation states are unstable to Heisenberg-type interaction terms. This implies that one can form a stable fixed-separation basis of zero-energy states, which can be used to encode some initial quantum information. This information is then preserved for a long period of time if the strength of the destabilizing perturbations is small. Interestingly, the Loschmidt echo for the staggered field perturbation shows an unusually strong revival at $t\approx 1,\lambda=1.0$, implying that one can recover the initial state even if the additional term generally destroys the coherence of the state.
@@ Acknowledgements.
Acknowledgments.— A. P., B. M., and M. S. were funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 853368). C. J. T. is supported by an EPSRC fellowship (Grant Ref. EP/W005743/1). The authors acknowledge the use of the UCL High Performance Computing Facilities (Myriad and Kathleen), and associated support services, in the completion of this work. R. M. and H. J. C. acknowledge funding from National Science Foundation Grant DMR 2046570 and the National High Magnetic Field Laboratory. The National High Magnetic Field Laboratory is supported by the National Science Foundation through NSF/DMR-1644779 and DMR-2128556 and the state of Florida.
@@ I.1 Character theory details
Let us start by explicitly writing the representations for the dihedral group $D_{2L}$, which is the anti-symmetry group $\mathcal{A}$ of our system (see Fig. 1 in the main text). It contains one-dimensional irreps labeled by $\kappa=0,L/2$ and $s=\pm 1$ with representation
@@
The characters are defined as traces of representation matrices, and hence it is easy to see that for 1D irreps $\chi_{\kappa,s}(\sigma^{\nu}\tau^{\mu})=s^{\nu}\omega_{\kappa}^{\mu}$, while for 2D irreps $\chi_{\kappa}(\sigma^{\nu}\tau^{\mu})=2\cos(2\pi\kappa\mu/L)$. Here, $\nu\in\{0,1\}$ and $\mu\in\{0,...,L-1\}$ denote the powers of the reflection and translation in the group element $g=\sigma^{\nu}\tau^{\mu}$. We can write this succinctly as $\chi_{\kappa,s}(\sigma^{\nu}\tau^{\mu})=d_{\kappa}s^{\nu}\cos(2\pi\kappa\mu/L)$, where $d_{\kappa}\in\{1,2\}$ is the dimensionality of irrep $\kappa$ and for 2D irreps we take $s=+1$.
Now, we turn to the discussion of characters for the internal symmetry of the model, SU(2). Consider the character function for the system of size $l$ (Hilbert space $\mathcal{H}_{l}$) from Eq. (4) in the main text. The trace can be expressed in some orthonormal product basis, for $\theta$ an element of the internal symmetry and $\tau$ generates a cycle of length $l$,
@@
[…]
$=\sum_{\alpha_{1}}\bra{\alpha_{1}}\rho_{\frac{1}{2}}(\theta)^{l}\ket{\alpha_{1}}=\mathrm{tr}[\rho_{\frac{1}{2}}(\theta)^{l}]\text{,}$ (S.3c)
where $\rho_{\frac{1}{2}}(\theta)$ is the spin-$\frac{1}{2}$ representation of SU(2) for a single site, and |$\ket{\alpha_{i}}$ is the basis for site $i$. The first step is to use $\tau$ to permute the product basis labels, then $\theta$ is split as a product operator acting as $\rho_{\frac{1}{2}}(\theta)$ on each site. The formula is then distributed into a product across sites, which is equivalent to taking the trace of a matrix power.
Now we can note that any spin-1/2 SU(2) matrix can be transformed into a diagonal form $\rho_{\frac{1}{2}}(\theta)=\operatorname{diag}(e^{i\theta},e^{-i\theta})$, giving us a straightforward way to compute the trace and characters,
@@
×$\quad\times\sum_{\mu=0}^{L-1}\sum_{\nu=0}^{1}d_{\kappa}s^{\nu}\cos\frac{2\pi\kappa\mu}{L}\prod_{l}(2\cos\theta l)^{c_{l}(g)}.$
This can be evaluated explicitly after first obtaining $l$-cycles and their counts $c_{l}(g)$, as follows. Two representative examples of how to calculate $\{l,c_{l}(g)\}$ are shown in Fig. S.1, with panel (a) presenting a translation-only case ($\nu=0$), and panel (b) showing a case with translation and reflection ($\nu=1$). For $\nu=0$, there are $c_{l}(g)=\text{gcd}(L,\mu)$ cycles of length $l=L/\text{gcd}(L,\mu)$. For $\nu=1$, there are $c_{2}(g)=L/2-1$ cycles of length $2$ and $c_{1}(g)=2$ cycles of length $1$.
Figure S.1: Diagram showing examples of $l$-cycles for a system of size $L=12$, for (a) $g=\sigma^{0}\tau^{8}$, and (b) $g=\sigma^{1}\tau^{3}$. Each color corresponds to one $l$-cycle. The corresponding cycle lengths $l$ and their counts $c_{l}(g)$ are listed in the tables below the panels.